OpenStudy (anonymous):

Andreas received a new book for his birthday. He read 1/8 of the book in the first week, then more pages in the second week. If he still has 162 pages to read, how many total pages are in his book?

OpenStudy (anonymous):

34 + 162 = 196 = 7/8 of number of pages in book therefore total pages = 196 * 8/7 = 224